OpenStudy (anonymous):

n(A)= 48, n(A')=40, n(B')=74 U= a universal set u= union. Find the number of elements in A u B. n(A u B)=_____________________________?

OpenStudy (mathmate):

Use \[A \cup A' = U\] and \[B \cup B' = U\] => n(U)=n(A)+n(A')=48+40=88 and n(U)=n(B)+n(B')=88 n(B)=88-n(B')=88-74=14 \[n(A \cup B) = n(A)+n(B)-n(A \cap B)\] \[=48+14-n(A \cap B)\] \[=62 - n(A \cap B)\] We cannot calculate a numerical value since the intersection of A and B is not known.